What is the vapor preesure of a solution made by dissolving 225 g of glucose in 358 mL of water at 30 c? The molar mass of glucose 180.2 g/mol. What is the vapor pressure lowering? The vapor pressure of pure water at 30 c is 31. 82 mmHg. Assure that the density of the solution is 1.00 g/mL


1
Expert's answer
2021-03-22T07:57:43-0400

Vapor pressure ,"P=Ps \u00d7 Ss"

"PsV= nRT"

225g of glucose= 225/180.2=1.25moles

Ps=(1.25×8.314×303.15)/3.58× 10-4

Ps=8800241pa